OpenStudy (anonymous):

find the cirsumfrenece of the circle given the circumfrence of a circle =3.14timeds d , where d =the diameter the radius =6.3

OpenStudy (anonymous):

d = 2 * r

OpenStudy (anonymous):

Take your radius, multiply it by 2, then substitute it into the formula for circumference.

OpenStudy (anonymous):

39.56

OpenStudy (anonymous):

39.564

OpenStudy (anonymous):

radius,r=6.3 diameter,d=2*6.3 circumference of a circle =3.14timesd=3.14*d circumference of a circle =3.14*2*6.3=39.564